OpenStudy (jojo12g):

1+5*3 ------ 2 A. 8 B. 9 C. 16 D. 18 its 9 right

OpenStudy (anonymous):

5x3=15+1=16. 16/2= 8. It is 8.

OpenStudy (anonymous):

a is the correct answer

OpenStudy (anonymous):

18

OpenStudy (jojo12g):

thank u so much

OpenStudy (anonymous):

np

OpenStudy (anonymous):

so its d bc 1+5=6*3=18